OpenStudy (anonymous):

How do you find the length of HI if GK=24,HJ=10, and GH=HI=IJ

OpenStudy (jchick):

Without the picture to go from, I'm assuming the points are in alphabetical order along the line. HJ = HI+IJ=2HI =>2HI=10 =>HI =IJ=GH =5 GK=HJ+JK =>JK=24-(GH+HI+IJ)=24-15=9 So the answers are 1. 5 2. 5 3. 5 4. 9 5. 5+5=10 6. 5+9=14